John is twice as old as Lubert. In 3 yrs, John will be 13 yrs less than twice as old as Lubert. Find their present ages. (Is this problem correct?)
Answer by ankor@dixie-net.com(22740)

John is twice as old as Lubert. In 3 yrs, John will be 13 yrs less than twice as old as Lubert. Find their present ages. (Is this problem correct?)
:
It is NOT correct, try to do it:
:
J is twice as old as L.
j = 2L
:
In three years, J will be 13 years less than twice as old as L.
j + 3 = 2(L+3) - 13
j + 3 = 2L + 6 - 13
j = 2L - 7 - 3
j = 2L - 10
:
Find their ages now.
replace j with 2L (from the 1st statement)
2L = 2L - 10
obviously, this is not possible.